OpenStudy (anonymous):

integrate {x^13 * (sin (x^7) dx}

OpenStudy (aravindg):

Did you try integration by parts?

OpenStudy (aravindg):

in btw : WELCOME TO OPENSTUDY!!

OpenStudy (anonymous):

Which one should be u and which one should be dv?

OpenStudy (anonymous):

Thanks

OpenStudy (callisto):

Do something before by parts. Let u=x^7 => du = 7x^6 dx => x^6 dx = du/7 \[\int x^{13}sin (x^7) dx\]\[=\int x^{7}sin (x^7) (x^6dx)\]\[=\int usin (u) (\frac{du}{7})\]\[=\frac{1}{7}\int usin(u)du\]Then, use integration by parts.

OpenStudy (anonymous):

Thank you Callisto

OpenStudy (callisto):

Welcome~ Do you understand how to get there?